A boclk slides down a frictionless incline with co
Answer #1

using the kinematic equations

V2 = u2 + 2 a s

and we have u = 0

3.802 = 2 X a X 6.80

14.44 = a X 13.6

a = 14.4 / 13.6

a = 1.05 m/s2

we have s = 1/2 a t2

3.4 = 0.5 X 1.05 X t2

t2 = 3.4/0.525

t2 = 6.47

t = 2.54 s

and to find the speed we have v = at

then v = 1.05 X 2.54

v = 2.667 m/s